# SpikySpace: A Spiking State Space Model for Energy-Efficient Time Series Forecasting

### v1

@@ IV-D Power-of-Two SiLU
It is easily seen that the value of $\bar{f}(x)$ is negligible when x<10x<-10. On the other hand, when $x>10$, the value of $\bar{f}(x)$ monotonically decreases and is bounded below by its asymptotic limit: $\lim_{x\rightarrow\infty}\bar{f}(x)\geq-0.229$. ∎
We provide the complete proof of Lemma 4 in the supplementary materials. Fig. 3 visually confirms the small deviations in both the function shapes and the derivatives of the two functions.
IV-E Training Process
As we adopted an ANN-to-SNN conversion method to obtain our SpikySpace, we first trained a quantized ANN model with PTsoftplus and PTSiLU as the activation functions and then converted it into an SNN model. Specifically, we followed BiT [34] and Sorbet [18] to quantize the activation into $n$ bits. The quantization is in a Learned Step-size Quantization(LSQ) manner. Represented as:
$X_{Q}=\alpha\cdot\operatorname{clip}\!\left(\operatorname{round}\!\left(\frac{X_{R}-\beta}{\alpha}\right),Q_{n},\;Q_{p}\right)+\beta$ (24)
We adopt a uniform quantization scheme with a learnable step size $\alpha$ and an optional offset $\beta$. According to [34], $\alpha$ can be initialized by
$\alpha^{*}=\frac{\bigl\lVert\mathbf{X}_{\mathrm{R}}\odot\mathbbm{1}_{\{\mathbf{X}_{\mathrm{R}}\geq 0.5\}}\bigr\rVert_{1}}{\,n_{\{\mathbf{X}_{\mathrm{R}}\geq 0.5\}}}$ (25)
and $\beta$ can be initialized as 0. Straight-through estimator (STE) is employed as commonly used in previous works to back-propagate the gradients to $\alpha$.
This quantization method covers both symmetric and asymmetric cases. Specifically, the real-valued tensor $X_{R}$ is first normalized by subtracting the offset $\beta$ and dividing by $\alpha$. The result is rounded to the nearest integer and clipped to the representable integer range $Q_{n},Q_{p}]$, which is determined by the bit width $b$. Finally, the quantized value is mapped back to the real domain as $X_{Q}$. When $\beta=0$ and $(Q_{n},Q_{p})=(-2^{b-1},\,2^{b-1}-1)$, the formulation reduces to symmetric quantization. When $\beta\neq 0$ and $(Q_{n},Q_{p})=(0,\,2^{b}-1)$, It becomes asymmetric quantization.
We note that our quantization method inherently retains certain floating-point operations for the scaling factor $\alpha$ and offset $\beta$. However, this does not pose a limitation, as prior studies [18] have shown that these parameters can be further quantized to power-of-two values, allowing multiplications to be replaced by efficient bit-shifting operations without noticeable accuracy degradation.
V Experimental Evaluation
TABLE I: Experimental results of time-series forecasting on four benchmarks with different prediction horizons ($L$). The best and second-best results in SNNs are highlighted in bold and underlined, respectively. $\uparrow$ ($\downarrow$) indicates that higher (lower) values are better.
Dataset Horizon GRU iTransformer SpikeTCN SpikeRNN iSpikformer SpikeSTAG SpikySpace
[…]
Metr-la 3 .803 .448 .864 .344 .845 .415 .784 .490 .805 .466 .873 .375 .894 .343
6 .761 .507 .849 .410 .799 .473 .731 .547 .765 .512 .822 .430 .848 .412
12 .682 .585 .763 .514 .718 .560 .661 .614 .723 .555 .734 .535 .760 .517
24 .614 .663 .538 .652 .602 .665 .557 .702 .549 .709 .590 .685 .610 .659
Avg. .715 .551 .754 .480 .741 .528 .683 .588 .711 .561 .755 .506 .778 .483
Pems-bay 3 .783 .479 .938 .269 .862 .401 .763 .527 .935 .276 .874 .384 .921 .304
6 .769 .504 .888 .362 .829 .448 .721 .571 .884 .369 .835 .439 .866 .395
12 .696 .638 .797 .488 .782 .504 .710 .582 .787 .499 .787 .537 .762 .528
24 .696 .638 .629 .659 .681 .582 .693 .599 .622 .665 .626 .661 .605 .683
Avg. .736 .565 .813 .444 .788 .484 .722 .570 .807 .452 .780 .505 .788 .478
Solar 3 .962 .508 .974 .562 .946 .205 .933 .246 .972 .217 .973 .246 .971 .176
6 .950 .548 .964 .584 .937 .252 .923 .278 .955 .218 .950 .272 .953 .223
12 .907 .569 .918 .575 .893 .409 .903 .343 .918 .295 .926 .315 .914 .301
24 .875 .572 .879 .541 .840 .541 .820 .425 .869 .372 .879 .390 .821 .434
Avg. .924 .549 .934 .566 .904 .352 .895 .323 .929 .275 .932 .306 .915 .284
Electricity 3 .983 .518 .983 .213 .974 .324 .984 .207 .982 .214 .987 .207 .994 .136
6 .981 .522 .977 .506 .970 .333 .978 .280 .974 .284 .986 .222 .992 .157
12 .980 .531 .977 .460 .968 .338 .979 .314 .973 .284 .985 .224 .990 .176
24 .972 .506 .977 .305 .963 .342 .964 .317 .974 .284 .984 .225 .990 .179
Avg. .979 .519 .978 .371 .969 .334 .976 .279 .976 .266 .985 .220 .992 .162
TABLE II: Statistics of the four benchmarks.
@@ V Experimental Evaluation
V-A Experimental Setups
V-A1 Dataset
We evaluate our model on four widely used multivariate time series forecasting benchmarks, spanning diverse application domains including traffic, solar energy, and electricity consumption. All datasets are divided chronologically into training, validation, and test sets without shuffling. The statistical details and split ratios are summarized in Table II.
  • METR-LA [10]: Traffic speed data collected from 207 loop detectors on Los Angeles highways, aggregated every 5 minutes. The dataset contains 34,272 samples, each with a sequence length of 12 and 207 variables. It is split into training, validation, and test sets with a ratio of (0.7, 0.2, 0.1)(0.7,\,0.2,\,0.1). This dataset captures strong spatial–temporal dependencies due to road network topology.
  • PEMS-BAY [10]: Another large-scale traffic dataset from the California Bay Area, sampled every 5 minutes from 325 sensors. It includes 52,116 samples, each of length 12 with 325 variables. The chronological split ratio is (0.7, 0.2, 0.1)(0.7,\,0.2,\,0.1). PEMS-BAY is widely adopted for benchmarking spatiotemporal forecasting models.
  • Solar-Energy [11]: Solar power production recorded from 137 photovoltaic plants across several years, exhibiting strong daily and seasonal periodicity. The dataset contains 52,560 hourly samples with 137 variables and a sequence length of 168. It is partitioned into training, validation, and test sets with a ratio of (0.6, 0.2, 0.2)(0.6,\,0.2,\,0.2).
  • Electricity [11]: Hourly electricity consumption data from 321 clients over multiple years, showing clear diurnal and weekly patterns. The dataset includes 26,304 samples, each with 321 variables and a sequence length of 168. It is divided into train/validation/test sets in the ratio (0.6, 0.2, 0.2)(0.6,\,0.2,\,0.2).
V-A2 Baselines
We compare SpikySPace with the following time series forecasting methods, including both ANN and SNN baselines:
  • GRU [35]: A recurrent neural network (RNN) model that captures temporal dependencies through gated recurrent units. It serves as a classical ANN baseline for sequence modeling.
  • iTransformer [36]: A state-of-the-art transformer variant that learns both instance-wise and channel-wise dependencies in multivariate time series, representing strong ANN performance under full-precision computation.
  • SpikeTCN [2]: A convolutional SNN that integrates temporal convolution with spike-based processing, highlighting the benefit of temporal feature extraction in event-driven computation.
  • SpikeRNN [2]: A recurrent SNN model that replaces analog activations with spiking neurons, enabling energy-efficient sequential modeling through temporal dynamics.
  • iSpikformer [2]: A spiking adaptation of the transformer architecture that replaces softmax attention and linear layers with spike-compatible modules, achieving competitive accuracy with reduced energy.
  • SpikeSTAG [29]: A hybrid spatio-temporal graph SNN designed for structured time-series data, combining graph connectivity with event-driven updates to model spatial correlations efficiently.
V-A3 Evaluation Metrics
@@ V-A4 Implementation Details
In all experiments, we set the batch size to 64 and the learning rate to $5\times 10^{-4}$. For each dataset, the forecasting horizons are set to 3, 6, 12, and 24 time steps. Our model is trained using the Adam optimizer with the mean squared error (MSE) as the loss function. During training, we set the maximum epoch number as 1000 while the early stop patience as 20. All experiments are conducted on a single NVIDIA A100 GPU with 80GB of memory. The SpikySpace model results we reported in Table I are using a timestep $T=3$, which means we use a spike train length as 3 to represent a scalar.
V-B Results
The evaluation results across four datasets and four prediction horizons for our models are shown in Table I. SpikySpace achieves the highest $R^{2}$ and lowest RRSE values on two datasets and near best on the other two, demonstrating strong generalization and stability across different domains and horizons. In particular, on the Electricity dataset, SpikySpace achieves $R^{2}=0.992$ and RRSE=0.162=0.162, surpassing both iSpikformer and iTransformer while using fully spiking computation. On traffic datasets like Metr-la, SpikySpace attains $R^{2}$ gains of $2.3\%$ respectively over the strongest SNN baselines, confirming that its selective scan mechanism effectively captures temporal dependencies under sparse updates. On the other two datasets, SpikySpace’s performance remains competitive with a smaller model size and less energy consumption. These results demonstrate that SpikySpace closes the accuracy gap between SNNs and dense ANNs while maintaining neuromorphic efficiency.
V-C Ablation Study
To better understand the contribution of each component in SpikySpace, we conducted a comprehensive ablation study focusing on the proposed activation approximations and the effect of timestep selection on forecasting performance.
V-C1 Approximated Operations
Figure 4: Ablation study results of PTSoftplus and PTSiLU on full-precision ANNs.
Bits PTsoftplus PTSiLU $R^{2}$ $\Delta R^{2}$
4 $\times$ $\times$ .607 N/A
$\checkmark$ $\times$ .611 +.004
$\times$ $\checkmark$ .611 +.004
$\checkmark$ $\checkmark$ .610 +.003
1 $\times$ $\times$ .612 N/A
$\checkmark$ $\times$ .605 -.007
$\times$ $\checkmark$ .616 +.004
$\checkmark$ $\checkmark$ .603 -.009
TABLE III: Ablation study on the impact of PTsoftplus and PTSiLU. ‘Bits’ in this table denotes the activation quantization bit-width.
To assess the effectiveness of the proposed PTsoftplus and PTSiLU activations, we conduct a controlled ablation against their full-precision counterparts. As shown in Fig. 4, replacing the standard Softplus and SiLU with our power-of-two variants achieves almost identical accuracy across all forecasting horizons, confirming that the approximations retain the expressive capacity of the original functions.
We then quantifies this effect under different activation bit-widths as shown in Table III. At 4-bit precision, both PTsoftplus and PTSiLU yield a small gain of $\Delta R^{2}=+0.004$, and their combination results in only +0.003+0.003 deviation from the Softplus–SiLU baseline. Even at 1-bit quantization, the accuracy drop remains below $1\%$ ($\Delta R^{2}=-0.009$), showing strong robustness to coarse quantization.
Overall, PTsoftplus and PTSiLU preserve smooth gating and differentiability while enabling efficient bit-shift implementations without costly exponentials or divisions.
V-C2 Timestep
TABLE IV: Ablation study results on timesteps. Tested across two datasets and all 4 horizons.
Dataset Horizon 1 3 7 15
[…]
Solar 3 0.971 0.176 0.971 0.176 0.970 0.179 0.969 0.181
6 0.952 0.225 0.952 0.224 0.949 0.232 0.946 0.239
12 0.911 0.306 0.913 0.302 0.905 0.317 0.904 0.318
24 0.833 0.420 0.821 0.434 0.830 0.424 0.838 0.413
Metr-la 3 0.894 0.344 0.894 0.343 0.894 0.343 0.895 0.343
6 0.847 0.413 0.848 0.412 0.847 0.413 0.847 0.413
12 0.757 0.520 0.760 0.517 0.759 0.518 0.758 0.519
24 0.604 0.664 0.610 0.659 0.608 0.661 0.613 0.656
In the main results shown in Table I, we adopted the T=3T{=}3 for the SpikySpace model, converted from the 2-bit activation ANNs. This setting provides a good balance between temporal resolution and computational efficiency in the selective scan.
To further analyze its impact, we conduct an ablation study by varying the internal timestep $T$ while keeping all other configurations identical and the results are shown in Table IV.
Overall, changing $T$ has little influence on forecasting accuracy across datasets and horizons, indicating that the discrete-time state evolution of SpikySpace is stable under different update granularities. However, when the timestep is reduced to T=1T{=}1, we observe a few undesired results on Metr-la at larger horizons. To avoid these rare but suboptimal behaviors, we adopt T=3T{=}3 for all reported experiments.
V-D Energy Analysis
TABLE V: Comparison of the model size, energy consumption, and performance across SpikySpace and baselines. The energy reduction is calculated relative to that of iTransformer. The measurements are based on the Electricity dataset with a prediction horizon of 3.
SpikySpace SSM 0.868 0.12 98.73% 0.994 0.136
SpikeSTAG LSTM, transformer 1.566 4.39 53.64% 0.987 0.207
iSpikformer transformer 1.634 3.19 66.30% 0.982 0.214
iTransformer transformer $\times$ 1.634 9.47 / 0.983 0.213
Unlike traditional deep networks, the energy consumption of an SNN is primarily determined by the time-window length $T$ and the spike rate $s$ [37]. In a conventional ANN, a neuron performs one MAC operation as $y\mathrel{+}=w\cdot x$. If we encode $x$ to spike train, which only consist 0s and 1s, we can replace the MACs with simpler ACCs and thus save energy. For example, $x=2/3$ can be encoded as $[1,1,0]$ with $T=3$. The two 1s indicates adding the weight twice. Mathematically, the ACC operations in SNN scales linearly with the total spike count, approximately $T\cdot s$ times.
In our spiking state-space models, as presented in Eq. 13, the transition matrix $\overline{\mathbf{A}}_{t}$ introduces a positive values and accumulates through time propagates, reaching the clamping maximum and leading to a spike rate toward $\sim\!100\%$. In this case, spike will occur at every time step and the spike train will consist of all 1s and no 0s. To address this, we scale the firing threshold as $\theta^{\prime}=T\cdot\theta$, so that according to the spiking algorithm showing in Algorithm 1, the $T$ spikes in the spike train collapses to a single spike, reducing the spike rate from $1$ to $1/T$. To preserve the overall signal magnitude, we also scale the weights by $T$ like $\textbf{W}^{\prime}=T\cdot\textbf{W}$, which keeps the output numerically consistent with the unscaled case.
With this notice, we reduce the average spike rate from $93\%$ to $26.8\%$, yielding an average energy of 0.12 mJ on the Electricity dataset with a horizon of 3.
The parameter size and energy consumption of our SpikySpace are shown in Table V. Compare with the state-of-the-art models, the parameter size of our model is about 53.1% to 55.4% while reduces estimated energy consumption by 98.73% and 96.24% compared to iTransformer and iSpikformer, respecitvely.
VI Discussion
VI-A Neuromorphic Compatibility
While current neuromorphic hardware platforms are still under active development and not yet widely accessible, we assess the hardware compatibility of SpikySpace by implementing the core computational operators within the Lava framework, targeting Intel’s Loihi architecture. These implementations verify that our proposed operations can be effectively mapped to existing neuromorphic substrates.
Although we map the model onto 45nm technology to optain the energy cost [38], these analyses do not fully capture hardware-level dynamics such as routing delay, synaptic buffering, or spike serialization overhead. Future work will focus on deploying SpikySpace on real neuromorphic platforms like Loihi [39] or BrainScaleS-2 [40] to obtain empirical latency and power measurements, validate our simulation assumptions, and enable further hardware-aware optimizations.
VI-B Forecasting Horizon
We observe that SpikySpace achieves superior accuracy at shorter forecasting horizons like 3, 6, and 12 steps, but the performance gain diminishes when the horizon came to 24 steps, especially for the Metr-la and Pems-bay datasets. We attribute this to the design of the state-space formulation. In multi-step forecasting, small prediction errors may accumulate across steps. Since SpikySpace updates its latent state through compact linear transitions without global re-contextualization, these accumulated deviations are not explicitly corrected.
Despite this limitation at longer horizons, strong performance on short forecasting horizons remains highly valuable in many real-world applications. In domains such as real-time traffic flow control, renewable energy regulation, and edge-based anomaly detection, the ability to deliver fast and energy-efficient short-term predictions is often more critical than maximizing long-horizon accuracy.
VII Conclusion
In this paper, we presented SpikySpace, the first fully spiking state space model that enables sparse, event-driven state updates by reformulating state-space dynamics into a spike-driven selective scan. Designed for energy-efficient time-series forecasting, SpikySpace extends the spiking paradigm beyond conventional recurrent or transformer structures, establishing a principled framework for neuromorphic sequence modeling.
To support spiking SSMs, we further proposed PTsoftplus and PTSiLU. These two functions serve as the counterparts of conventional Softplus and SiLU activations, achieving both computational efficiency and biological plausibility. Evaluated on four benchmarks of time seires forecasting, SpikySpace achieved performance comparable to state-of-the-art models with substantial energy savings.
AI-Generated Content Acknowledgment
This manuscript made limited use of ChatGPT (GPT-5 2025) solely for language polishing to enhance readability and grammar. The tool was not used in drafting or revising any scientific content, including conceptual development, technical descriptions, data analysis, figures, or conclusions. All scientific content was developed and verified by the authors.

### v2

@@ C.4 Proof of Lemma 4
It is easily seen that $\bar{f}(x)$ is negligible when x<10x<-10. For $x>10$, the function value $\bar{f}(x)$ eventually decreases to the limit $\lim_{x\rightarrow\infty}\bar{f}(x)\geq-0.229$. ∎
Appendix D Detailed Experimental Settings
D.1 Datasets
In this section, we provide a detailed description of the dataset used in our experiments. All datasets are divided chronologically into training, validation, and test sets without shuffling. The statistical details and split ratios are summarized in Table 5.
Table 5: Statistics of the four benchmarks.
@@ D.1 Datasets
  • METR-LA (Li et al., 2018): Traffic speed data collected from 207 loop detectors on Los Angeles highways, aggregated every 5 minutes. The dataset contains 34,272 samples, each with a sequence length of 12 and 207 variables. It is split into training, validation, and test sets with a ratio of (0.7, 0.2, 0.1)(0.7,\,0.2,\,0.1). This dataset captures strong spatial–temporal dependencies due to road network topology.
  • PEMS-BAY (Li et al., 2018): Another large-scale traffic dataset from the California Bay Area, sampled every 5 minutes from 325 sensors. It includes 52,116 samples, each of length 12 with 325 variables. The chronological split ratio is (0.7, 0.2, 0.1)(0.7,\,0.2,\,0.1). PEMS-BAY is widely adopted for benchmarking spatiotemporal forecasting models.
  • Solar-Energy (Lai et al., 2018): Solar power production recorded from 137 photovoltaic plants across several years, exhibiting strong daily and seasonal periodicity. The dataset contains 52,560 hourly samples with 137 variables and a sequence length of 168. It is partitioned into training, validation, and test sets with a ratio of (0.6, 0.2, 0.2)(0.6,\,0.2,\,0.2).
  • Electricity (Lai et al., 2018): Hourly electricity consumption data from 321 clients over multiple years, showing clear diurnal and weekly patterns. The dataset includes 26,304 samples, each with 321 variables and a sequence length of 168. It is divided into train/validation/test sets in the ratio (0.6, 0.2, 0.2)(0.6,\,0.2,\,0.2).
D.2 Baselines
We compare SpikySpace with the following methods:
  • GRU (Cho et al., 2014): A recurrent neural network (RNN) model that captures temporal dependencies through gated recurrent units. It serves as a classical ANN baseline for sequence modeling.
  • iTransformer (Liu et al., ): A state-of-the-art transformer variant that learns both instance-wise and channel-wise dependencies in multivariate time series, representing strong ANN performance under full-precision computation.
  • SpikeTCN (Lv et al., 2024): A convolutional SNN that integrates temporal convolution with spike-based processing, highlighting the benefit of temporal feature extraction in event-driven computation.
  • SpikeRNN (Lv et al., 2024): A recurrent SNN model that replaces analog activations with spiking neurons, enabling energy-efficient sequential modeling through temporal dynamics.
  • iSpikformer (Lv et al., 2024): A spiking adaptation of the transformer architecture that replaces softmax attention and linear layers with spike-compatible modules, achieving competitive accuracy with reduced energy.
  • SpikeSTAG (Hu et al., 2025): A hybrid spatio-temporal graph SNN designed for structured time-series data, combining graph connectivity with event-driven updates to model spatial correlations efficiently.
D.3 Evaluation Metrics
@@ D.4 Implementation Details
In all experiments, we set the batch size to 64 and the learning rate to $5\times 10^{-4}$. For each dataset, the forecasting horizons are set to 3, 6, 12, and 24 time steps. Our model is trained using the Adam optimizer with the mean squared error (MSE) as the loss function. During training, we set the maximum epoch number as 1000 while the early stop patience as 20. All experiments are conducted on a single NVIDIA A100 GPU with 80GB of memory. The SpikySpace model results we reported in Table 1 are using a timestep $T=3$, which means we use a spike train length of 3 to represent a scalar.
D.5 Quantization Details
The straight-through estimator (STE) is employed as commonly used in previous works to back-propagate the gradients to $\alpha$.
Our quantization method covers both symmetric and asymmetric cases. Specifically, the real-valued tensor $X_{R}$ is first normalized by subtracting the offset $\beta$ and dividing by $\alpha$. The result is rounded to the nearest integer and clipped to the representable integer range $Q_{n},Q_{p}]$, which is determined by the bit width $b$. Finally, the quantized value is mapped back to the real domain as $X_{Q}$. When $\beta=0$ and $(Q_{n},Q_{p})=(-2^{b-1},\,2^{b-1}-1)$, the formulation reduces to symmetric quantization. When $\beta\neq 0$ and $(Q_{n},Q_{p})=(0,\,2^{b}-1)$, It becomes asymmetric quantization.